# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

### v3

@@ 4. Purely elliptic type
Our goal in this section is to prove that tall critical points of $g$ modulo $\Phi$ that have purely elliptic type satisfy the hypotheses of Proposition 3.4 (see Proposition 4.5 below). An important ingredient is the local classification of such points (see Lemma 4.4).
We begin by giving an important example and describing the symplectic form on (complexity one) local models.
@@
Remark 4.2.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a complexity one local model, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho:H\to(S^{1})^{h+1}$ with homogeneous moment map $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$. Let $i^{*}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{*}$ be the projection induced by the inclusion $H\hookrightarrow T$. Equip $T^{*}T\times{\mathbb{C}}^{h+1}$ with the symplectic form given by the sum of the pullbacks of the canonical symplectic form on $T^{*}T$ and the standard symplectic form on ${\mathbb{C}}^{h+1}$. The $H$-action on $T^{*}T\times{\mathbb{C}}^{h+1}\cong T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $h\cdot(t,\beta,w):=(ht,\beta,\rho(t^{-1})w)$ has a moment map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ that takes $(t,\beta,z)$ to $i^{*}(\beta)-\Phi_{H}(z)$. The reduced space at 0 is smooth, and the map $Y\to(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ taking $t,\alpha,z]$ to $t,\alpha+\Phi_{H}(z),z]$ is a diffeomorphism. By definition, the symplectic form $\omega_{Y}$ on $Y$ is the pullback of the reduced symplectic form along this diffeomorphism.
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in M$ be a critical point of $g$ modulo $\Phi$ that has purely elliptic type. Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be the local model for $p$, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\to(S^{1})^{h+1}$. Given $b\in{\mathbb{R}}^{h+1}\smallsetminus\rho_{*}({\mathfrak{h}})$, a neighborhood of $T\cdot p$ in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ is isomorphic to a neighborhood of $T\cdot[1,0,0]$ in $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f_{b}=(\Phi_{Y},g_{b}))$, where $g_{b}\colon Y\to{\mathbb{R}}$ is given by $g_{b}([t,\alpha,z])=\sum_{i=0}^{h}b_{i}|z_{i}|^{2}$.
@@
is a symplectomorphism between the local model $Y$ and the reduced space $(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ discussed in Remark 4.2.
@@
Since $p$ has tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$, where $\rho\colon H\to(S^{1})^{h+1}$ is the symplectic slice representation at $p$. Hence, by Lemma 4.4 and the Morse Lemma, after possibly replacing $g$ by g-g, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, that $g([t,\alpha,z])=|z|^{2}$, and that $p=[1,0,0]$. Let $N$ be the degree of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$. By Lemma 2.7, the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism; moreover, it restricts to a diffeomorphism from $({Y/\!/T})\smallsetminus\{[t,0,0]\}$ to ${\mathbb{C}}\smallsetminus\{0\}$. By Lemma 2.9, there exists a positive real number $C$ such that $g([t,\alpha,z])=C|P(z)|^{\frac{2}{N}}$ for all $t,\alpha,z]\in\Phi_{Y}^{-1}(0)$. Let $S\colon{\mathbb{C}}\to{\mathbb{C}}$ be the homeomorphism that, in polar coordinates, takes $(r,\theta)$ to $(\sqrt{C}\,r^{\frac{1}{N}},\theta)$; it restricts to a diffeomorphism ${\mathbb{C}}\smallsetminus\{0\}\to{\mathbb{C}}\smallsetminus\{0\}$. Define $\Psi\colon{Y/\!/T}\to{\mathbb{C}}\simeq{\mathbb{R}}^{2}$ by $\Psi([t,\alpha,z])=S(P(z))$. Since
@@
We observe that, by Lemma 2.8, if $N=1$ the map $\Psi$ constructed in the above proof is a diffeomorphism.
@@ 5. The semitoric case
Let $Y={\mathbb{C}}^{2}$, let $\omega_{Y}$ be the standard symplectic form, and set $\Phi_{Y}(z_{1},z_{2}):=\frac{1}{2}\left(|z_{1}|^{2}-|z_{2}|^{2}\right)$ and $g_{Y}(z_{1},z_{2}):={\mathfrak{I}}(z_{1}z_{2})={\mathfrak{I}}(P(z_{1},z_{2}))$, where $P\colon Y\to{\mathbb{C}}$ is the defining polynomial for the $S^{1}$-action induced by $\Phi_{Y}$ (cf. Example 2.6).
@@
Otherwise, since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ either has purely elliptic type or is focus-focus. By Propositions 3.4 and 4.5, and Lemma 5.4, M//T{M/\!/T} can be given the structure of a smooth closed, connected, oriented surface such that $\overline{g}:{M/\!/T}\to{\mathbb{R}}$ is a Morse function with no critical point of index 1. By Proposition 5.6, the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Since $T$ is connected, it follows that $f^{-1}(0,c)$ is also connected. ∎
@@ 6. Non-degenerate singular points
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$. By Lemma 3.2, $p$ is a critical point of $g$ modulo $\Phi$. Hence, the $(T\times{\mathbb{R}})$-orbit $\mathcal{O}$ through $p$ coincides with the $T$-orbit $\mathfrak{O}$ through $p$ and the Taylor polynomial $T^{1}_{0}(j^{*}g)$ of degree one at $0\in{\mathbb{C}}^{h+1}$ vanishes. Thus we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}=(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $T^{2}_{0}(j^{*}f)=(T^{2}_{0}(j^{*}\Phi),T^{2}_{0}(j^{*}g))=(\Phi_{H},T^{2}_{0}(j^{*}g))$, where $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ is the homogeneous moment map for the $H$-action.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is non-degenerate and tall; let $N$ be the degree of the defining polynomial of the local model for $p$. Then $p$ has at most one non-elliptic block; furthermore,
@@
Let $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ be the defining polynomial of $Y$. Depending on the degree $N$ of $P$, we may assume that we are in one of the following four cases (cf. Examples 2.5 and 2.6).
@@
By Lemma 2.8, the inclusion $i\colon{\mathbb{C}}\to Y$ sending $z_{0}$ to $1,0,(z_{0},0,\ldots,0)]$ induces a diffeomorphism $\overline{i}\colon{\mathbb{C}}\to{Y/\!/T}$. By definition of the inclusion $j$ and by (6.4),
@@
Finally, assume that $p$ has a hyperbolic block. Then 0 is a non-degenerate critical point of $i^{*}g$ of index 1. Hence, by the equivariant Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 0 that is an equivariant diffeomorphism onto an open set such that $(i\circ\psi)^{*}g(x,y)=xy$ for all $(x,y)\in V$. Let $\pi\colon{\mathbb{R}}^{2}/{\mathbb{Z}}_{2}\to{\mathbb{R}}^{2}$ be the map that sends $[x,y]$ to $(xy,x^{2}-y^{2})$; it is a homeomorphism onto ${\mathbb{R}}^{2}$ and a diffeomorphism on the complement on $(0,0)$. Setting $U:=(\overline{i}\circ\overline{\psi})(V/{\mathbb{Z}}_{2})$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\pi\circ(\overline{i}\circ\overline{\psi})^{-1},$ the result follows.
@@
where the signs in the first term are independent. Hence, since $p$ is non-degenerate, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}\neq 0$. If $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, then $p$ has one focus-focus block, and the remaining blocks are elliptic; otherwise $p$ has purely elliptic type. It follows from (6.6), and Lemmas 2.7 and 2.9, that the defining polynomial identifies the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ with the map $\mathbb{C}$$\to{\mathbb{R}}$ taking $w\in{\mathbb{C}}$ to $a_{x}{\mathfrak{R}}(w)+a_{y}{\mathfrak{I}}(w)+2a_{\rho}|w|$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is homeomorphic to $\mathbb{R}$} exactly if $p$ has a focus-focus block, and is a single point if $p$ has purely elliptic type.
@@
Since $p$ is non-degenerate, the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=0,\dots,h\}$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]$ to $T^{2}_{0}(j^{*}g)$ is the point $[t,0,0]$. Moreover, since all the eigenvalues of the symplectic endomorphism corresponding to any element of the above span are imaginary, $p$ has purely elliptic type by Remark 6.1. ∎
@@
Let $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi_{Y},g))$ be an integrable system, where $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a tall local model. Assume that $p=[1,0,0]$ is a non-degenerate critical point of $g$ modulo $\Phi$ with $g(p)=0$. If $j\colon{\mathbb{C}}^{h+1}\to Y$ is the inclusion $z\mapsto[1,0,z]$, then

### v4

@@ 4. Purely elliptic type
Our goal in this section is to prove that tall critical points of $g$ modulo $\Phi$ that have purely elliptic type satisfy the hypotheses of Proposition 3.6 (see Proposition 4.4 below). An important ingredient is the local classification of such points (see Lemma 4.3). We begin by giving an important example.
@@
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in M$ be a critical point of $g$ modulo $\Phi$ that has purely elliptic type. Let $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ be the local model for $p$, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\to(S^{1})^{h+1}$. Given $b\in{\mathbb{R}}^{h+1}\smallsetminus\rho_{*}({\mathfrak{h}})$, a neighborhood of $T\cdot p$ in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ is isomorphic to a neighborhood of $T\cdot[1,0,0]$ in $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f_{b}=(\Phi_{Y},g_{b}))$, where $g_{b}\colon Y\to{\mathbb{R}}$ is given by $g_{b}([t,\alpha,z])=\sum_{i=0}^{h}|z_{i}|^{2}b_{i}$.
@@
is a symplectomorphism between the local model $Y$ and the reduced space $(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$; see Remark 2.2.
@@
Since $p$ has tall local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, Remark 2.5 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$, where $\rho\colon H\to(S^{1})^{h+1}$ is the symplectic slice representation at $p$. Hence, by Lemma 4.3 and the Morse Lemma, after possibly replacing $g$ by g-g, we may assume that $M$ is a local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, that $\Phi=\Phi_{Y}$, that $g([t,\alpha,z])=|z|^{2}$, and that $p=[1,0,0]$. Let $N$ be the degree of the defining monomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$. By Lemma 2.10, the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism; moreover, it restricts to a diffeomorphism from $({Y/\!/T})\smallsetminus\{[1,0,0]\}$ to ${\mathbb{C}}\smallsetminus\{0\}$. By Lemma 2.12, there exists a positive real number $C$ such that $g([t,\alpha,z])=C|P(z)|^{\frac{2}{N}}$ for all $t,\alpha,z]\in\Phi_{Y}^{-1}(0)$. Let $S\colon{\mathbb{C}}\to{\mathbb{C}}$ be the homeomorphism that, in polar coordinates, takes $(r,\theta)$ to $(\sqrt{C}\,r^{\frac{1}{N}},\theta)$; it restricts to a diffeomorphism ${\mathbb{C}}\smallsetminus\{0\}\to{\mathbb{C}}\smallsetminus\{0\}$. Define $\Psi\colon{Y/\!/T}\to{\mathbb{C}}\simeq{\mathbb{R}}^{2}$ by $\Psi([t,\alpha,z])=S(P(z))$. Since
@@
We observe that, by Lemma 2.10 (and 2.9), if $N=1$ the map $\Psi$ constructed in the above proof is a diffeomorphism.
@@ 5. The semitoric case
Let $Y={\mathbb{C}}^{2}$, let $\omega_{Y}$ be the standard symplectic form, and set $\Phi_{Y}(z_{1},z_{2}):=\frac{1}{2}\left(|z_{1}|^{2}-|z_{2}|^{2}\right)$ and $g_{Y}(z_{1},z_{2}):={\mathfrak{I}}(z_{1}z_{2})={\mathfrak{I}}(P(z_{1},z_{2}))$, where $P\colon Y\to{\mathbb{C}}$ is the defining monomial for the $S^{1}$-action induced by $\Phi_{Y}$ (cf. Example 2.8).
@@
Otherwise, since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ either has purely elliptic type or is focus-focus. By Propositions 3.6 and 4.4, and Lemma 5.4, M//T{M/\!/T} can be given the structure of a smooth closed, connected, oriented surface such that $\overline{g}:{M/\!/T}\to{\mathbb{R}}$ is a Morse function with no critical point of index 1. By Proposition 5.6, the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Since $T$ is connected, it follows that $f^{-1}(0,c)$ is also connected. ∎
@@ 6. Non-degenerate singular points
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$. By Lemma 3.4, $p$ is a critical point of $g$ modulo $\Phi$. Hence, the $(T\times{\mathbb{R}})$-orbit $\mathcal{O}$ through $p$ coincides with the $T$-orbit $\mathfrak{O}$ through $p$ and the Taylor polynomial $T^{1}_{0}(j^{*}g)$ of degree one at $0\in{\mathbb{C}}^{h+1}$ vanishes. Thus we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}=(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $T^{2}_{0}(j^{*}f)=(T^{2}_{0}(j^{*}\Phi),T^{2}_{0}(j^{*}g))=(\Phi_{H},T^{2}_{0}(j^{*}g))$, where $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ is the homogeneous moment map for the $H$-action.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is non-degenerate and tall; let $N$ be the degree of the defining monomial of the local model for $p$. Then $p$ has at most one non-elliptic block; furthermore,
@@
Let $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ be the defining monomial of $Y$. Depending on the degree $N$ of $P$, we may assume that we are in one of the following four cases (cf. Examples 2.7 and 2.8).
@@
By Lemma 2.10, $\overline{P}$ is a diffeomorphism; moreover, the inclusion $i\colon{\mathbb{C}}\to Y$ sending $z_{0}$ to $1,0,(z_{0},0,\ldots,0)]$ induces the inverse diffeomorphism $\overline{i}\colon{\mathbb{C}}\to{Y/\!/T}$. By definition of the inclusion $j$ and by (6.4),
@@
Finally, assume that $p$ has a hyperbolic block. Then 0 is a non-degenerate critical point of $i^{*}g$ of index 1. Hence, by the equivariant Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}\simeq{\mathbb{C}}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 0 that is a ${\mathbb{Z}}_{2}$-equivariant diffeomorphism onto an open set such that $(i\circ\psi)^{*}g(x,y)=xy$ for all $(x,y)\in V$. Let $\pi\colon{\mathbb{R}}^{2}/{\mathbb{Z}}_{2}\to{\mathbb{R}}^{2}$ be the map that sends $[x,y]$ to $(xy,x^{2}-y^{2})$; it is a homeomorphism onto ${\mathbb{R}}^{2}$ and a diffeomorphism on the complement on $(0,0)$. Setting $U:=(\overline{i}\circ\overline{\psi})(V/{\mathbb{Z}}_{2})$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\pi\circ(\overline{i}\circ\overline{\psi})^{-1},$ the result follows.
@@
where the signs in the first term are independent. Hence, since $p$ is non-degenerate, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}\neq 0$. If $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, then $p$ has one focus-focus block, and the remaining blocks are elliptic; otherwise $p$ has purely elliptic type. It follows from (6.6), and Lemmas 2.10 and 2.12, that the defining monomial identifies the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ with the map $\mathbb{C}$$\to{\mathbb{R}}$ taking $w\in{\mathbb{C}}$ to $a_{x}{\mathfrak{R}}(w)+a_{y}{\mathfrak{I}}(w)+2a_{\rho}|w|$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is homeomorphic to $\mathbb{R}$} exactly if $p$ has a focus-focus block, and is a single point if $p$ has purely elliptic type.
@@
Since $p$ is non-degenerate, the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=0,\dots,h\}$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]$ to $T^{2}_{0}(j^{*}g)$ is the point $[1,0,0]$. Moreover, since all the eigenvalues of the symplectic endomorphism corresponding to any element of the above span are imaginary, $p$ has purely elliptic type by Remark 6.1. ∎
@@
Let $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi_{Y},g))$ be an integrable system, where $Y=T\times_{H}({\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1})$ is a tall local model. Assume that $p=[1,0,0]$ is a non-degenerate critical point of $g$ modulo $\Phi$ with $g(p)=0$. If $j\colon{\mathbb{C}}^{h+1}\to Y$ is the inclusion $z\mapsto[1,0,z]$, then